OpenStudy (anonymous):

Sam borrowed $2875 from David at a simple interest rate of 3.5% p.a. Determine the time period of the loan in days, if Sam paid $2924.65 to pay off the loan

OpenStudy (demonchild99):

is there any more info to give???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Sorry, that other stuff was wrong. here is what I'm pretty sure you're supposed to do: calculate simple interest using the formula I=PRT, where I = interest, P = principal, R = the interest rate, and T = time. so they gave us P (2875), R (3.5%, which as a decimal that is written as .035). and we can find I by subtracting P from the given final amount (2924.65-2875=49.65, so 49.65 is our I).... so plug those in and solve for T

OpenStudy (anonymous):

like this: \[49.65=(2875)\times(.035)\times(T)\]

OpenStudy (anonymous):

then multiply the two numbers on the right and divide by that number on both sides.

OpenStudy (anonymous):

that number meaning the answer to the multiplication problem

OpenStudy (anonymous):

So \[2875\times.035=100.625\] Take that number and divide it into 49.65 \[\frac{ 49.65 }{ 100.625 }=.493416149\] The answer is T...